OpenStudy (lastdaywork):

Need a second opinion - If f is an even function defined on the interval (-5,5), then find four real values of x satisfying the equation -

OpenStudy (lastdaywork):

\[f(x) = f \left( \frac{ x+1 }{ x+2 } \right)\]

OpenStudy (lastdaywork):

My answer - \[\frac{ -1\pm \sqrt 5 }{ 2 } , \frac{ -3\pm \sqrt 5 }{ 2 }\]

OpenStudy (lastdaywork):

Answer in the book - \[\frac{ -3\pm \sqrt 5 }{ 2 } , \frac{ +3\pm \sqrt 5 }{ 2 }\]

OpenStudy (anonymous):

Is that an even function?

OpenStudy (lastdaywork):

Yes, f is an even function.

OpenStudy (anonymous):

I thought "even" meant f(-x) = f(x), or have I gotten it wrong?

OpenStudy (nincompoop):

you got it right, doug

OpenStudy (lastdaywork):

@douglaswinslowcooper f satisfies f(-x) = f(x) Now we need to find four values of x satisfying f(x) = f((x+1)/(x+2))

ganeshie8 (ganeshie8):

how did u get the other two values ?

OpenStudy (lastdaywork):

I used - \[x = \frac{ x+1 }{ x+2 }\]

ganeshie8 (ganeshie8):

thats incorrect right ? whenever (x, f(x)) is a point on graph, then (-x, f(x)) will be a point also... if x1 and x2 output the same value, then x1 = -x2

ganeshie8 (ganeshie8):

f(x1) = f(x2) => x1 = -x2

ganeshie8 (ganeshie8):

im just messing wid the symmetry..

ganeshie8 (ganeshie8):

correct right, f(2) = f(-2) when f is even 2 = --2

OpenStudy (lastdaywork):

If f is an even function; cant we say f(x) = f(y) => x=±y ??

ganeshie8 (ganeshie8):

oh in this particular problem yes.. LastDayWork.. pls throw some light on how u got the other two values...

ganeshie8 (ganeshie8):

hmm no i dont think so : cuz, if f(2) = f(3) then 2 =/= +- 3

OpenStudy (lastdaywork):

No, I mean for an even function f If f(x) = f(y) then x = y or x = -y

OpenStudy (nincompoop):

?

ganeshie8 (ganeshie8):